# OpenVLThinker: An Early Exploration to Complex Vision-Language Reasoning via Iterative Self-Improvement

### v1

@@ 5 Conclusion
In this work, we introduced OpenVLThinker-7B, a LVLM enhanced through an iterative self-improving process combining SFT and RL to enable complex CoT reasoning. Our results demonstrate that integrating R1-style reasoning into LVLMs effectively boosts their multimodal reasoning performance across benchmarks. This initial result highlights the promising potential for future exploration of complex reasoning in vision-language contexts.
@@ Appendix A Additional Empirical Study
Does Complex Reasoning Matter for VQA?
We additionally investigated whether complex, multi-step reasoning provides significant performance gains over standard (non-R1) reasoning in visual tasks. In this study, we use the ConTextual [52] validation set of 100 VQA examples, aiming to disentangle the roles of image grounding and textual reasoning. As similar to our first distillation process, we separately employ a vision-language model for caption generation and a pure-text model for reasoning. The image description generated by the captioning model is then fed into one of two text-based models: DeepSeek-R1-Distill-14B (an R1-style reasoner) or Qwen2.5-14B-Instruct (a standard instruction-tuned model). This setup allows us to isolate the impact of R1 reasoning from the effects of the underlying vision encoder.
@@
(1) R1 reasoning outperforms standard methods.
When provided with identical captioned inputs, DeepSeek-R1-Distill-14B achieves higher accuracy than Qwen2.5-14B-Instruct. Moreover, its performance can match (or even surpass) the direct QA accuracy of its own captioning model (LLaVA-v1.6-34B), despite potential information loss from translating the image into text.
(2) Sampling benefits complex reasoners.
Increasing the number of sampled reasoning chains ($k=2$ or $k=4$) leads to larger performance gains for R1 models than for standard Qwen models, indicating that the multi-step reasoning approach can more effectively converge on correct solutions when multiple hypotheses are explored.
(3) Image grounding quality matters.
We observe that richer and more precise captions significantly enhance final VQA accuracy. When captions are more detailed (e.g., from GPT-4o), the improvements from complex reasoning are especially pronounced.
@@
Caption Refinement via Feedback: Limited Effectiveness.
We also investigated whether a one-round feedback loop could improve the quality of captions and thus final VQA performance. Concretely, DeepSeek-R1-Distill-14B was prompted to list missing or ambiguous details in the initial captions generated by LLaVA-v1.6-34B. The captioning model then re-generated a “refined” description incorporating this feedback. Table 5 shows that the refined captions did not produce major accuracy improvements, suggesting that a single feedback pass is insufficient for significantly enriching image descriptions.
@@ Appendix B Experiment Details
We thank LLaMA-Factory [^4] https://github.com/hiyouga/LLaMA-Factory and EasyR1555https://github.com/hiyouga/EasyR1 for open-sourcing the training framework that we used for SFT and GRPO. In Table 7 and 8, we detail the hyperparameters that we used for SFT, GRPO and inference. We further lay out the prompts we used for generating image captions.
@@
Figure 8: The image to the visual reasoning task that we showed in section 4.2
Figure 9: The image to the demonstrated reasoning example in section 1. Question: What is the length of side XY? Choices: (A) 4 (B) 5 (C) 6 (D) 9.
Figure 10: The reasoning length distribution of DeepSeek-R1-Distill-Qwen-14B given the image descriptions and questions from VQA datasets.
Figure 11: Pass@4 accuracy of DeepSeek-R1-Distill-14B, which takes image descriptions and questions as inputs, on the six data sources we selected for SFT-Iter1.
@@ Appendix C Output Examples
Repetitive reasoning example after SFT without filtering But the question is asking for the route and endpoint. So, the route would be the highway number, and the endpoint would be the destination. Wait, but the signs are arranged vertically. So, the top signs are the routes, and the bottom ones are the destinations. So, if you’re at the junction and you’re facing the signs, you can see both. So, the route is the highway number, and the endpoint is the destination. But the question is phrased as "the route and endpoint reached by continuing in a straight line." So, if you’re at the junction and you’re facing the signs, you can see both the routes and the destinations. So, the route would be the highway number, and the endpoint would be the destination. Wait, but the instruction is a bit ambiguous. It says "continuing in a straight line," which might mean following the route indicated by the sign. So, if you’re at the junction and you’re facing the signs, you can see both the routes and the destinations. So, the route is the highway number, and the endpoint is the destination. But the question is asking for the route and endpoint. So, the route would be the highway number, and the endpoint would be the destination. Wait, but the signs are arranged vertically. So, the top signs are the routes, and the bottom ones are the destinations. So, if you’re at the junction and you’re facing the signs, you can see both. So, the route is the highway number, and the endpoint is the destination. But the question is phrased as "the route and endpoint reached by continuing in a straight line." So, if you’re at the junction and you’re facing the signs, you can see both the routes and the destinations. So, the route would be the highway number, and the endpoint would be the destination.
Model Output at GRPO-Iter1 Gradient Step 3 user If the parallelogram is cut along the dashed line, what is the length of the resulting rectangle’s longer side? assistant <think> Okay, so I’ve got this image of a parallelogram with some measurements. The image says the top base is 21 cm, the bottom base is 17 cm, and the height is 8 cm labeled at the bottom base. There’s a dashed line inside the parallelogram from the top right corner to the central point on the left base, making a right angle with the left base. </think> The dashed line is a median, which splits the parallelogram into two congruent triangles. When the parallelogram is cut along this line, each resulting trapezoid has a longer side of 21 cm, as the median connects opposite sides. Since the dotted line is perpendicular, it forms a rectangle along the base, making the longer side 21 cm. <answer>21 cm</answer> ground truth 21 cm
Model Output at GRPO-Iter1 Gradient Step 5 user How many non-cube objects are there after removing all brown cubes? assistant <think> Okay, so I need to figure out how many non-cube objects there are after removing all brown cubes. Let me start by un[0/1833]ing the image provided. The image has five objects on a plain, light surface. They are: 1. A small, shiny golden cube on the left. 2. A larger, metallic green sphere near the center-left. 3. A smaller, brown cube near the center-right. 4. Another green sphere, slightly smaller than the metallic one, on the right. 5. A teal, cylindrical object in the foreground. First, the question is about non-cube objects after removing all brown cubes. So, I need to identify which objects are not cubes initially and then subtract the brown cube since it’s being removed. The cubes are the small golden one and the smaller brown one. Removing the brown cube means we’re taking it out of the count. Now, the non-cube objects are: - The large metallic green sphere. - The smaller green sphere. - The teal cylinder. So, after removing the brown cube, we’re left with three objects. Therefore, the number of non-cube objects is three. </think> The image contains five objects: one small golden cube, one large green metallic sphere, one small brown cube, another small green sphere, and a teal cylinder. The question asks how many non-cube objects remain after removing all brown cubes. Initially, the non-cube objects are the two spheres (one green metallic and one green) and the teal cylinder. The two cubes (golden and brown) are removed, leaving the three non-cube objects. <answer>3</answer> ground truth 3
Model Output at GRPO-Iter1 Gradient Step 45 user Does Forest Green have the highest peak value? assistant <think> Okay, so I need to figure out if Forest Green has the highest peak value in the graph described. Let me start by understanding the information given. The graph has seven lines, each with different colors: Orange Red, Coral, Forest Green, Tomato, Red, Aqua, and Crimson. The y-axis ranges from about 68 to 69. The x-axis goes from 0 to 100, in intervals of 20. Looking at the image, each line starts at a point on the y-axis and then either increases, decreases, or fluctuates. The task is to determine if Forest Green has the highest peak. First, I’ll go through each line one by one to see their starting points and trends. 1. **Orange Red**: Starts around 68 and decreases gradually. So, its peak is at the start, which is around 68. 2. **Coral**: Starts around 68 and decreases slightly. So, at x=0, it’s also around 68. 3. **Forest Green**: Starts at 69 and increases slightly. So, its peak is higher than the others mentioned so far. 4. **Tomato**: Starts at 68 and increases gradually. So, it peaks around 69, same as Forest Green. 5. **Red**: Starts at 68 and increases sharply, reaching above 69. So, Red is one of the highest. 6. **Aqua**: Starts at 68 and increases slightly. So, around 69. 7. **Crimson**: Starts at 68 and increases slightly. Same as Aqua. </think> The Forest Green line starts at 69 and increases slightly, meaning its peak value is 69. However, the Red line starts at 68 and increases sharply to a value exceeding 69. Therefore, Red has the highest peak value. <answer>No</answer> ground truth No

### v2

@@ 6 Conclusion
In this work, we proposed a new perspective on LLM reasoning as actions at inference time, signified by keywords such as “wait”. We thus interpret the roles of SFT as action highlighting that efficiently surfaces desired actions by distilling a reasoning model’s demonstrations. On the other hand, RL makes improvement on basis provided by SFT. Based on this intuition, we introduced OpenVLThinker-7B, a LVLM enhanced through an iterative self-improving process combining SFT and RL to enable complex CoT reasoning. Our results demonstrate that integrating R1-style reasoning into LVLMs effectively boosts their multimodal reasoning performance across benchmarks. With only three SFT-RL cycles and 12K training examples, the model raises average accuracy on six diverse visual-reasoning benchmarks to 46.6 %, with a 2 % absolute gain over its base model and on par with proprietary systems such as GPT-4o.
Limitations. Our experiments span six established benchmarks, yet they do not exhaustively probe robustness in other tasks or real-world settings. In addition, we validated the method only on a 7B model as a proof of concept. While the approach should scale to larger backbones (e.g., 32B) and likely yield further gains, such exploration requires substantially greater computational resources.
@@ Appendix A Additional Empirical Study
Does Complex Reasoning Matter for VQA? We additionally investigated whether complex, multi-step reasoning provides significant performance gains over standard (non-R1) reasoning in visual tasks. In this study, we use the ConTextual [65] validation set of 100 VQA examples, aiming to disentangle the roles of image grounding and textual reasoning. As similar to our first distillation process, we separately employ a vision-language model for caption generation and a pure-text model for reasoning. The image description generated by the captioning model is then fed into one of two text-based models: DeepSeek-R1-Distill-14B (an R1-style reasoner) or Qwen2.5-14B-Instruct (a standard instruction-tuned model). This setup allows us to isolate the impact of R1 reasoning from the effects of the underlying vision encoder.
@@
(1) R1 reasoning outperforms standard methods. When provided with identical captioned inputs, DeepSeek-R1-Distill-14B achieves higher accuracy than Qwen2.5-14B-Instruct. Moreover, its performance can match (or even surpass) the direct QA accuracy of its own captioning model (LLaVA-v1.6-34B), despite potential information loss from translating the image into text.
(2) Sampling benefits complex reasoners. Increasing the number of sampled reasoning chains ($k=2$ or $k=4$) leads to larger performance gains for R1 models than for standard Qwen models, indicating that the multi-step reasoning approach can more effectively converge on correct solutions when multiple hypotheses are explored.
(3) Image grounding quality matters. We observe that richer and more precise captions significantly enhance final VQA accuracy. When captions are more detailed (e.g., from GPT-4o), the improvements from complex reasoning are especially pronounced.
@@
Caption Refinement via Feedback: Limited Effectiveness. We also investigated whether a one-round feedback loop could improve the quality of captions and thus final VQA performance. Concretely, DeepSeek-R1-Distill-14B was prompted to list missing or ambiguous details in the initial captions generated by LLaVA-v1.6-34B. The captioning model then re-generated a “refined” description incorporating this feedback. Table 5 shows that the refined captions did not produce major accuracy improvements, suggesting that a single feedback pass is insufficient for significantly enriching image descriptions.
@@ Appendix B Experiment Details
We thank LLaMA-Factory [^6] https://github.com/hiyouga/LLaMA-Factory and EasyR1777https://github.com/hiyouga/EasyR1 for open-sourcing the training framework that we used for SFT and GRPO. In Table 7 and 8, we detail the hyperparameters that we used for SFT, GRPO and inference. We further lay out the prompts we used for generating image captions. Experiments were conducted on GPU clusters to the similar level of NVIDIA H100 80GB GPU. SFT/Distillation requires 30 minutes and RL requires 20 hours for each iteration. In addition, distillation data generation with verification requires about 8 hours.
@@ Appendix C Output Examples
Repetitive reasoning example after SFT without filtering But the question is asking for the route and endpoint. So, the route would be the highway number, and the endpoint would be the destination. Wait, but the signs are arranged vertically. So, the top signs are the routes, and the bottom ones are the destinations. So, if you’re at the junction and you’re facing the signs, you can see both. So, the route is the highway number, and the endpoint is the destination. But the question is phrased as "the route and endpoint reached by continuing in a straight line." So, if you’re at the junction and you’re facing the signs, you can see both the routes and the destinations. So, the route would be the highway number, and the endpoint would be the destination. Wait, but the instruction is a bit ambiguous. It says "continuing in a straight line," which might mean following the route indicated by the sign. So, if you’re at the junction and you’re facing the signs, you can see both the routes and the destinations. So, the route is the highway number, and the endpoint is the destination. But the question is asking for the route and endpoint. So, the route would be the highway number, and the endpoint would be the destination. Wait, but the signs are arranged vertically. …
Figure 12: The image to the demonstrated reasoning example in section 1.
Figure 13: The image to the visual reasoning task that we show for SFT-Iter3 and GRPO-Iter3 output below.
Figure 14: The image to the task that we show for SFT-Iter3 and GRPO-Iter3 output below.
Example of SFT-Iter3 incorrect output and GRPO-Iter3 Correct output to the same question Question: The derivative of f(x) at x=2 is ____ that at x=5 Choices: (A) larger than (B) equal to (C) smaller than SFT-Iter3
<think>
Okay, let me see. The question is asking about the derivative of the function f(x) = |2x - 3| + 1 at x=2 compared to x=5. First, I need to recall what the derivative represents. The derivative gives the slope of the tangent line at a point on the function. So I need to look at the slopes at those two points. </think> The function ( f(x) = |2x - 3| + 1 ) is an absolute value function, which means it has a V-shape. The vertex of this V-shape occurs where the expression inside the absolute value is zero, i.e., ( 2x - 3 = 0 ), which gives ( x = $\frac{3}{2}$ ). The derivative of ( f(x) ) will be zero at the vertex because the function changes direction there. For ( $x<\frac{3}{2}$ ), the function is decreasing, and for ( $x>\frac{3}{2}$ ), it is increasing. The slopes on either side of the vertex are negative and positive, respectively. At ( x = 2 ) (which is to the right of the vertex), the function is increasing, so the derivative is positive. At ( x = 5 ) (which is further to the right), the function is also increasing, so the derivative is positive but steeper because the function is increasing faster. Since the function is increasing faster at ( x = 5 ) than at ( x = 2 ), the derivative at ( x = 5 ) is larger. <answer>(C)</answer> ================== GRPO-Iter3 <think> Okay, let’s see. The question is asking about the derivative of the function f(x) = |2x - 3| + 1 at x=2 compared to x=5. First, I need to recall what the derivative represents. The derivative at a point gives the slope of the tangent line at that point. So I need to look at the slopes at x=2 and x=5. </think> The function ( f(x) = |2x - 3| + 1 ) is an absolute value function shifted up by 1. The absolute value function ( |2x - 3| ) has a V-shape with a vertex at ( x = $\frac{3}{2}$ ). The derivative of ( |2x - 3| ) is ( 2 ) for ( x $>\frac{3}{2}$ ) and ( -2 ) for ( x < $\frac{3}{2}$ ). Adding 1 to the function doesn’t affect the derivative’s magnitude but shifts the y-values up. At ( x = 2 ), which is greater than ( $\frac{3}{2}$ ), the derivative is ( 2 ). At ( x = 5 ), which is also greater than ( $\frac{3}{2}$ ), the derivative is also ( 2 ). Since the derivative is the same at both points, the slopes are equal. <answer>(B)</answer> Ground truth: (B)
Example of SFT-Iter3 incorrect output and GRPO-Iter3 Correct output to the same question Question: Find $m\angle 3$. Choices: (A) 28 (B) 38 (C) 52 (D) 62 SFT-Iter3
<think>
Okay, let me see. The problem is asking for the measure of angle 3 in the given quadrilateral. The hint suggests looking at the angles provided, which are 52 degrees and 38 degrees. Hmm, the quadrilateral is a rectangle, so all angles should be 90 degrees. But the hint is confusing me. Let me check the options again. The angles 52 and 38 are given, but they’re not the ones I need to find. The question is about angle 3, which is opposite angle 1. Since opposite angles in a rectangle are equal, angle 3 should be the same as angle 1. But angle 1 is not given directly. Let me see, the hint is saying to look at the angles provided. The 52 and 38 are supplementary angles, which means they add up to 180 degrees. But angle 3 is not supplementary to 52 or 38. It’s the opposite angle, so it should be the same as angle 1. Since angle 1 is not given, I need to use the options. The only angle that could be the same as angle 1 is the one that’s not given directly. The options are A, B, C, D. The only one that’s not supplementary to 52 or 38 is 62. So angle 3 must be 62 degrees. </think> The problem states that the quadrilateral is a rectangle, and all angles in a rectangle are 90 degrees. The hint provided the angles 52° and 38°, which are supplementary angles (52° + 38° = 90°). However, angle 3 is the opposite angle to angle 1, which means they are equal because opposite angles in a rectangle are congruent. Since angle 1 is not given directly, we need to look at the options provided. The only angle that fits the properties of opposite angles in a rectangle is the one that is not supplementary to the given angles. The options are A (28), B (38), C (52), and D (62). Since angle 1 is not given, the only angle that could be the same as angle 3 is the one that is not supplementary to 52° or 38°. The only angle that fits this criterion is 62°, as it is the only one not listed as supplementary to the given angles. <answer>(D)</answer> ================== GRPO-Iter3 <think> Okay, let’s see. The question is asking for the measure of angle 3. The image provided is a rectangle with some angles labeled. The hint suggests looking at the angles given, which are 52 degrees and 38 degrees. Hmm, the rectangle has right angles, so each angle should be 90 degrees. The angles 52 and 38 are supplementary because they add up to 90 degrees (90 - 52 = 38). But angle 3 is not one of those. It’s the angle opposite the 38-degree angle, so it should be the same. Wait, the question is about angle 3, which is the one opposite the 38-degree angle. So it should be 38 degrees too, right? </think> The problem states that angle 3 is opposite the 38-degree angle. In a rectangle, opposite angles are equal. Since angle 3 is opposite the 38-degree angle, angle 3 must also be 38 degrees. <answer>(B)</answer> Ground truth: (B)